Q.#2 Please solve the following equations. a) 31-5x) et Final Answer b)-1(x2 - 1)2 = 2 In V2x2 – 3x+1 Final Answer
Answer #1

Here, Here, (a) = amen. from 07. 38x38. Tiem, the general formula, axalaman. 6-5x + 2x² Note! when, x shen y. a = a., then x=.. = 6 * osť xo! Here x = 3 orn 6 I en (x-1) = 2 en 2x²-3n+1 0 From General formula, enx = y.enx) Brantatt - Here, en c2²-1)²x²-1 = 2x² – 3x + 1 2x²-x² -3x+1+1=0 Pa x²-3x+2=0 x= - b ± √b²-490 2a = + 3 ± √9 - (4x1x2) a-1 b=-3 coa = 3 + √9-8 - 311 = 40